write and equation of the line that passes through (4, -3) and is perpendicular to the line y = 4x -1

Answer by mahikab(11)

When figuring out the perpendicular line, we will first have the original slop (of 4) and multiply it by -1 (now -4) and write its reciprocal (-1/4). Now our line is y = -1/4x + b.
Now we input our coordinate of (4,-3) in for the corresponding values.
The x-coordinate is 4.
The y-coordinate is -3.
Our y = -1/4x + b now becomes:
-3 = -1/4(4) + b
-3 = -1 + b
b = -3 + 1
b = -2
So your equation is y = -1/4x - 2.